4. The following are matters of standard deviation. Choose four numbers from a integer between 1 and 10 including 1 and 10 according to the following conditions. Repetition is allowed at this time. You can choice the number already selected. a. Choose four integer values to be the smallest standard deviation. b. Choose four integer values to make the biggest standard deviation. c. Can I combine multiple times in a or b? xplain.
media%2Fb7d%2Fb7d6cde3-30f7-42bd-95ab-27
Answer #1

6)

a)

245, ơ 40 We need to compute Pr( 200). The corresponding z-value needed to be computed is X-μ 200-245 40 =-1.125 Therefore, we get that Pr(X > 200) = Pr( z 200 245 40 40Pr(Z 1.125) 1-0.13030.8697 The following is obtained graphically: Normal Distribution: Prix> 200)-0.8697 0.0100 0.0090 0.0080 0.0070 0.0060 0.0050 0.0040 0.0030 0.0020 0, 0010 0, 0000 100 150 200 250 300 350 400 450

b)

X = \mu - Z_{\alpha }* \sigma

X = 245 - -0.5244* 40 = 224.024